OpenStudy (anonymous):

Find the Difference. Express in simplest form. Z+3/ x+2 - (-6z-5)/ z+2

OpenStudy (alexwee123):

\[z+\frac{ 3 }{ x }+2-\frac{ \left( -6z-5 \right) }{ z }+2\]

OpenStudy (alexwee123):

is that the equation? ^^